OpenStudy (anonymous):

A chemist wishes to make 3L of a 7% acid solution by mixing a 9% and a 4% acid solution. How many litres of each solution should the chemist use? My answer is 2.4L for the 4% solution and 600mL for the 9% solution. I used the allegation method. However when I tried to double check my answers, retracing it, the numbers became whacky. Anyone has an idea why?

OpenStudy (anonymous):

that was wrong i would solve \[.09x+.04(3-x)=3\times .07\] for \(x\) but @FoolForMath is the master or allegation frankly i had never heard of it before he mentioned it

OpenStudy (anonymous):

this method is easy enough though, start with \[9x+4(3-x)=21\] and you will be done in a trice

OpenStudy (anonymous):

I think allegation is easier, lol. Yours looks complicated with all the symbols. Tic, tac, toe!

OpenStudy (anonymous):

what? it is a simple linear equation, like from elementary algebra

OpenStudy (anonymous):

but if you think allegation is easier, then be my guest

OpenStudy (anonymous):

Elementary is primary school right?

OpenStudy (anonymous):

can also solve \[7x+3y=21\] \[x+y=3\] but it is clearly the same

OpenStudy (anonymous):

primary school if you go to a real school US public school, might see it in 7th or 8th grade inner city school you will see it when you take remedial classes at the local community college

OpenStudy (anonymous):

that can't be right since total is 3

OpenStudy (anonymous):

that's what I did! Than it turned downfall with the division, let me see if I can get a pic

OpenStudy (anonymous):

your total is 7!

OpenStudy (anonymous):

4.2+2.8=7L

OpenStudy (anonymous):

A chemist wishes to make 3L of a 7% acid

OpenStudy (anonymous):

No, no it's 7% acid solution, 3L of the solution

OpenStudy (anonymous):

right so answer cannot be 4.2 liters of one and 2.8 liters of the other

OpenStudy (anonymous):

i am sure it is superfast, but i don't know how to do it. i am sure @FoolForMath can adjust, but the answer is not that

OpenStudy (anonymous):

I got an answer that makes sense 2400ml and 600 ml = 3L, but I can't get the same results when I double check it.

OpenStudy (anonymous):

\[9x+4(3-x)=21\] \[9x+12-3x=21\] \[6x+12=21\] \[6x=9\] \[x=1.5\]

OpenStudy (anonymous):

1.5 liters of 9% and the rest (1.5 also) liters of 4%

OpenStudy (anonymous):

damn i messed up

OpenStudy (anonymous):

Huh? I think I take the mess up Trophy here. Huh?

OpenStudy (anonymous):

I multiplied with 7 in the last step, it should be 3.

OpenStudy (anonymous):

\[9x+4(3-x)=21\] \[9x+12-4x=21\] \[5x=9\] \[x=1.8\] there! helps to know the distributive law

OpenStudy (anonymous):

so my answers wrong than?

OpenStudy (anonymous):

Let me do this again,  9% ----> \(\frac 3 5 \times 3 =1.8 L\) 4% ----> \(\frac 2 5 \times 3 = 1.2 L\)